Expected Return Expected Return Expected Return
Probability A B C
Best state 0.25 40% 25% 15%
Good state 0.25 30% 20% 12%
Normal state 0.25 20% 15% 9%
Poor state 0.25 10% 10% 6%
beta - 1.8 1.1 0.7

Complete the following table.

A B C
Expected average return (e.g., 10.00%)   %   %   %
Standard deviation (e.g., 10.00%)   %   %   %

If a portfolio consists of A, B, and C is structured as follows, complete the table.

A B C
Portfolio weight 0.20 0.30 0.50
Portfolio
Expected average return (e.g., 10.00%)   %
Answer #1

Expected average return: average return = probability & Enpected return A = 0.25x40% + 0.25x 30%. + 0.25x 20% +0.25*10% = 101Standard deviation - P(-) where, p= probability x = return of security 7 = average de mean return À I return (A) (A-A) (A - Asecurity B Tum (8) -B (6-8) 6. 45 1. 5 56-25 0.22 0 25625 0-25 । -2.5 6.2 07 10 -1.5 56.4 P (1-3) 14.663 1.562 J-5625 [५.6625of Expected average return portfolio = wyr + Wa xr2 + Wz XY3 where w,= weights of individual returns = average return already